# PoCo: Point Context Cluster for RGBD Indoor Place Recognition

### v1

@@ Abstract
We present a novel end-to-end algorithm (PoCo) for the indoor RGB-D place recognition task, aimed at identifying the most likely match for a given query frame within a reference database. The task presents inherent challenges attributed to the constrained field of view and limited range of perception sensors. We propose a new network architecture, which generalizes the recent Context of Clusters (CoCs) to extract global descriptors directly from the noisy point clouds through end-to-end learning. Moreover, we develop the architecture by integrating both color and geometric modalities into the point features to enhance the global descriptor representation. We conducted evaluations on public datasets ScanNet-PR and ARKit with 807 and 5047 scenarios, respectively. PoCo achieves SOTA performance: on ScanNet-PR, we achieve R@1 of $64.63\%$, a $5.7\%$ improvement from the best-published result CGis ($61.12\%$); on Arkit, we achieve R@1 of $45.12\%$, a $13.3\%$ improvement from the best-published result CGis ($39.82\%$). In addition, PoCo shows higher efficiency than CGis in inference time ($1.75$X-faster), and we demonstrate the effectiveness of PoCo in recognizing places within a real-world laboratory environment.
@@ I Introduction
For mobile robots, localization is a critical capability that enables them to determine their position within a known environment [ijcai2021p603]. Visual Place Recognition (VPR) is typically formulated as an image retrieval problem, i.e., retrieving candidate frames from a database by comparing similarities to a given query frame [lowry2015visual, ijcai2021p603]. Consequently, the design of an accurate place recognition system plays an important role in localization tasks by matching current observations with previously visited scenarios from the database [zhu2023r2former, ming2022cgis, arandjelovic2016netvlad]. It finds extensive use in applications like navigation [stumm2013probabilistic, mirowski2018learning], and loop closure in SLAM [bresson2017simultaneous, lowry2015visual] etc. However, the place recognition problem is not trivial. There could be various challenging environments [lowry2015visual, ijcai2021p603, liu2023survey, zhu2023r2former, ming2022cgis] with illumination changes, objects changes, dynamic objects, etc. Furthermore, different sensors could be employed in specific environmental conditions. For example, Lidar [uy2018pointnetvlad] and Mono-camera [zhu2023r2former, lowry2015visual] are used for outdoor place recognition, while RGB-D cameras [ming2022cgis, yudin2022hpointloc] are widely used for indoor scenarios. How to effectively process the perception data from the sensor is also challenging [qi2017pointnet, cocs, wu2023pointconvformer]. In this paper, our main focus is on addressing the problem of indoor RGB-D point cloud-based place recognition.
@@
RGB-D Place Recognition is Under Development: A common approach for place recognition involves extracting global descriptors from both query and database frames, followed by ranking retrieved candidate frames based on the computed similarities of these global descriptors [lowry2015visual, ming2022cgis, yudin2022hpointloc]. Many approaches have been proposed for RGB place recognition [arandjelovic2016netvlad, zhu2023r2former, lowry2015visual], which are mostly for outdoor place recognition because RGB camera captures rich, long-range and dense color information. For indoor scenarios, other than the RGB camera, the RGB-D camera shows more capability of perceiving dense depth images, whereas the geometric information should also be considered for the place recognition task. PointNet-VLAD [uy2018pointnetvlad], MinkLoc-3D [komorowski2021minkloc3d] and Indoor DH3D [du2020dh3d] process point clouds for place recognition, but they are not designed to process color information. To utilize both color and geometric information, Sizikova et al. [inbook] process RGB and depth images separately by convolutional layers and concatenate them as a joint descriptor. To completely fuse the geometric and color information in the global descriptors, CGiS-Net [ming2022cgis] utilizes KP-Conv [thomas2019kpconv] to process both color and geometric data. In our approach, we present a novel structure to improve the efficacy of processing geometric information and also propose a better feature encoder for RGB-D place recognition tasks.
@@ II Related Works
Visual Place Recognition: Place recognition is often formulated as an image retrieval problem, where the candidate frames are ranked by calculating similarities between their global descriptors and the descriptor of a query frame [ming2022cgis, zhu2023r2former]. Traditional place recognition methods extract classical RGB image features, SIFT, SURF, etc [david2004distinctive, cummins2008fab, galvez2012bags] and use the features to compose descriptors, e.g. Bag-of-Words [sivic2008efficient], to match the frames. Then learning-based approaches genrate more representative and imperical features for visual place recognition and improved the performance [arandjelovic2016netvlad, lowry2015visual]. Especially, CNN-based methods [uy2018pointnetvlad, jin2017learned, gordo2017end] become predominant methods for visual place recognition tasks. By extending feature-engineered VLAD into a data-driven learning approach, NetVLAD-based approaches [arandjelovic2016netvlad] have resulted in higher accuracy in terms of matching than the traditional methods. However, extracting 3D spatial geometric features from static images is still challenging. On the other hand, numerous works directly consume point clouds as their inputs.  [uy2018pointnetvlad] employs PointNet [qi2017pointnet] structure and upgrade  [arandjelovic2016netvlad] to directly extract features from a 3d point cloud. To improve the limitation of  [uy2018pointnetvlad] on capturing local geometric structures,  [komorowski2021minkloc3d] proposed to compute more discriminative 3D descriptors based on a sparse voxelized point cloud representation and sparse 3D convolutions.  [du2020dh3d] designed a Siamese network that jointly learns 3D local feature detection and description directly from raw 3D points. While point clouds offer numerous advantages such as encoded 3D spatial information and viewpoint invariance, it is arguable that integrating 3D features with 2D color features into the descriptors could potentially enhance performance further as this integration exploits the complementary nature of 3D and 2D features, leveraging the strengths of both modalities for more comprehensive scene representation.
@@
Features Processing: Traditional methods [rublee2011orb, cummins2008fab] use manual-crated features for place recognition. After that CNN-based methods demonstrated better performance in different vision tasks [uy2018pointnetvlad, jin2017learned, gordo2017end] and also place recognition. NetVLAD-based approaches [arandjelovic2016netvlad] show good performance for RGB place recognition, and KPConv-based approaches [Shi2023LiDARBasedPR, ming2022cgis] show promising results in point cloud place recognition. Vision transformers (ViTs) [dosovitskiy2020image] is the next generation of the feature extractor and outperform CNN-based approaches in multiple vision tasks [touvron2021training, berton2022deep, zhu2023r2former]. Recently, a new light-weighted feature extraction method Context of Clusters (CoCs) [cocs], is proposed and demonstrates state-of-the-art performance in different vision tasks but is more computationally effective than ViTs. However, CoCs is only designed for RGB features. In our work, we generalize it to RGB-D point cloud feature extraction.
@@ III-B Architecture of Our PoCo Model
Reducer Blocks: Each reducer block is used to (1) downsample points population from the last layer; (2) estimate relative geometric relationships and aggregate embedding features for reduced points. For points downsampling, we choose the Farthest Point Sampling (FPS) strategy [eldar1997farthest] for its excellent computational efficiency and ability to preserve the structure of the point cloud. Then for each point $p\in\mathbf{P}^{l+1}_{r}$, we apply KNN [kramer2013k] to choose the K nearest points from $\mathbf{P}^{l}_{r}$ as shown in Fig. 3, where $\mathbf{P}^{l}_{r}$ and $\mathbf{P}^{l+1}_{r}$ represent input and output point sets in the reducer block. The feature aggregation from $\mathbf{P}^{l}_{r}$ to $\mathbf{P}^{l+1}_{r}$ is based on PointConvFormer [wu2023pointconvformer] as shown in Eq. 1, where $\mathcal{N}(p)$ represents $p$’s neighbor set. $f_{1,2,3,4}(\cdot)$ are learnable Linear Layers, and $\psi(\cdot)$ is a multi-head attention module calculating the similarity score between $f(p_{k})$ and $f(p)$.
@@
Aggregation: For the point set $\mathbf{P}^{l+1}_{r}$, we downsample it to generate center points $\mathbf{P}^{l+1}_{c}$, where the cardinality $|\mathbf{P}^{l+1}_{r}|=N_{r}>|\mathbf{P}^{l+1}_{c}|=N_{c}$. Each center feature $\mathbf{f}_{c}$ is initialized as the mean vector of its K spatially nearest point features $\{\mathbf{f}_{p}\}$ from $\mathbf{P}^{l+1}_{r}$. As is shown in Fig. 4,in the aggregation stage, we firstly calculate the cosine similarity scores between all the centers and all the points: $s_{c,p}=\sigma(\alpha\cdot cosine(\mathbf{f}_{p},\mathbf{f}_{c})+\beta)$, where $\sigma(\cdot)$ is the sigmoid activation function and $\alpha$ and $\beta$ are trainable parameters. Then point features are aggregated into center points according to previous computed similarities. The center feature $\mathbf{f}_{c}$ is incorporated as an anchor feature for numerical stability as well as further emphasize the locality in Eq. 3
$\mathbf{f}_{c}=\frac{1}{\mathcal{C}}\left(\mathbf{f}_{c}+\sum_{p=1}^{N_{r}}\mathbf{s}_{c,p}\cdot\mathbf{f}_{p}\right),$ (3)
@@
Dispatch: The previous learned similarities are used to dispatch the aggregated center features to their member cluster points via Eq.4:
@@ III-C Training
The circle loss, $\mathcal{L}_{c}$, is defined as Equation 7. $m=0.2$ is a margin to reinforce the optimization. $cl(\cdot)$ is a clamp function, which clamps the value to 0 if it is negative. $\gamma=1$ is a hyperparameter. $s_{p}$ and $s_{n}$ are cosine similarities of the query-positive frame pair and query-negative frame pair, respectively. The circle loss guides the similarities between the query and positive cases to 1 and negative cases to 0.
@@
$\mathcal{L}_{c}=softplus(\log\sum\exp(\gamma a_{n}(s_{n}-\delta_{n}))-$ (6)
@@
Triplet loss, $\mathcal{L}_{t}$, is defined as Equation 8. Triplet loss is to maximize the difference between the distances of the query-positive frame pair and query-negative frame pair. The distance between the query and database descriptors is calculated by the L2 distance function $d(\cdot,\cdot)$. The margin is set as $m=0.2$.
@@
In the final loss function, we combine the two losses $\mathcal{L}=\alpha\mathcal{L}_{c}+\beta\mathcal{L}_{t}$. In experiments, we find the best weights for these two terms as $\alpha=10$ and $\beta=0.1$ via hyper-parameters tuning. By assigning a higher weight to triplet loss, the model can rapidly converge to a specific range, facilitating the differentiation between positive and negative examples. Subsequently, circle loss aids in refining the estimation process, resulting in more accurate similarity values that closely align with the boundaries.
@@ IV-A Dataset and Training Settings
Our PoCo model is implemented by Pytoch and trained on 8 Tesla-V100 GPUs. The input points are constrained to 2000 points through voxelization. Besides $\left\{r,g,b\right\}$ and $\left\{x,y,z\right\}$, we also use the normal vector $\left\{n_{x},n_{y},n_{z}\right\}$ of each point to provide additional relative geometric information as in Eq. 2, which is estimated by the plane within a 0.2-meter radius of the point. For the training strategy, we use cosine annealing scheduler [loshchilov2016sgdr] with Adam optimizer [kingma2014adam]. The learning rate is from $10^{-4}$ to $10^{-7}$.
Query Frames Vanilla-RGB-PR PoCo Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Query Frames PoCo w/o Color PoCo Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@ IV-B Implementations of Comparison and Ablation Study
To evaluate the performance of our approach, PoCo, we compare with different SOTA indoor RGB-D and point cloud-based place recognition approaches, CGiS-Net [ming2022cgis], MinkLoc-3D [komorowski2021minkloc3d], NetVLAD [arandjelovic2016netvlad], PointNet-VLAD [uy2018pointnetvlad], Indoor DH3D [du2020dh3d]. All these approaches are trained and tested in the two datasets, separately.

### v2

@@ Abstract
We present a novel end-to-end algorithm (PoCo) for the indoor RGB-D place recognition task, aimed at identifying the most likely match for a given query frame within a reference database. The task presents inherent challenges attributed to the constrained field of view and limited range of perception sensors. We propose a new network architecture, which generalizes the recent Context of Clusters (CoCs) to extract global descriptors directly from the noisy point clouds through end-to-end learning. Moreover, we develop the architecture by integrating both color and geometric modalities into the point features to enhance the global descriptor representation. We conducted evaluations on public datasets ScanNet-PR and ARKit with 807 and 5047 scenarios, respectively. PoCo achieves SOTA performance: on ScanNet-PR, we achieve R@1 of $64.63\%$, a $5.7\%$ improvement from the best-published result CGis ($61.12\%$); on Arkit, we achieve R@1 of $45.12\%$, a $13.3\%$ improvement from the best-published result CGis ($39.82\%$). In addition, PoCo shows higher efficiency than CGis in inference time ($1.75$X-faster), and we demonstrate the effectiveness of PoCo in recognizing places within a real-world laboratory environment. Video: https://youtu.be/D8dObAeMiCw;
@@ I Introduction
For mobile robots, localization is a critical capability that enables them to determine their position within a known environment [ijcai2021p603]. Visual Place Recognition (VPR) is typically formulated as an image retrieval problem, i.e., retrieving candidate frames from a database by comparing similarities to a given query frame [lowry2015visual, ijcai2021p603]. Consequently, the design of an accurate place recognition system plays an important role in localization tasks by matching current observations with previously visited scenarios from the database [zhu2023r2former, ming2022cgis, arandjelovic2016netvlad]. It finds extensive use in applications like navigation [stumm2013probabilistic, mirowski2018learning], and loop closure in SLAM [lowry2015visual] etc. However, the place recognition problem is not trivial. There could be various challenging environments [lowry2015visual, ijcai2021p603, liu2023survey, zhu2023r2former, ming2022cgis] with illumination changes, objects changes, dynamic objects, etc. Furthermore, different sensors could be employed in specific environmental conditions. For example, Lidar [uy2018pointnetvlad] and Mono-camera [zhu2023r2former, lowry2015visual] are used for outdoor place recognition, while RGB-D cameras [ming2022cgis, yudin2022hpointloc] are widely used for indoor scenarios. How to effectively process the perception data from the sensor is also challenging [qi2017pointnet, cocs, wu2023pointconvformer]. In this paper, our main focus is on addressing the problem of indoor RGB-D point cloud-based place recognition.
@@
RGB-D Place Recognition is Under Development: A common approach for place recognition involves extracting global descriptors from both query and database frames, followed by ranking retrieved candidate frames based on the computed similarities of these global descriptors [lowry2015visual, ming2022cgis, yudin2022hpointloc]. Many approaches have been proposed for RGB place recognition [arandjelovic2016netvlad, zhu2023r2former, lowry2015visual], which are mostly for outdoor place recognition because RGB camera captures rich, long-range and dense color information. For indoor scenarios, other than the RGB camera, the RGB-D camera shows more capability of perceiving dense depth images, whereas the geometric information should also be considered for the place recognition task. PointNet-VLAD [uy2018pointnetvlad], MinkLoc-3D [komorowski2021minkloc3d] and Indoor DH3D [du2020dh3d, yang2022fd] process point clouds for place recognition, but they are not designed to process color information. To utilize both color and geometric information, Sizikova et al. [inbook] process RGB and depth images separately by convolutional layers and concatenate them as a joint descriptor. To completely fuse the geometric and color information in the global descriptors, CGiS-Net [ming2022cgis] utilizes KP-Conv [thomas2019kpconv] to process both color and geometric data. In our approach, we present a novel structure to improve the efficacy of processing geometric information and also propose a better feature encoder for RGB-D place recognition tasks.
@@ II Related Works
Visual Place Recognition: Place recognition is often formulated as an image retrieval problem, where the candidate frames are ranked by calculating similarities between their global descriptors and the descriptor of a query frame [ming2022cgis, zhu2023r2former]. Traditional place recognition methods extract classical RGB image features, SIFT, SURF, etc [david2004distinctive, cummins2008fab, galvez2012bags] and use the features to compose descriptors, e.g. Bag-of-Words [sivic2008efficient], to match the frames. Then learning-based approaches generate more representative and emperical features for visual place recognition and improved the performance [arandjelovic2016netvlad, lowry2015visual]. Especially, CNN-based methods [uy2018pointnetvlad, jin2017learned, gordo2017end] become predominant methods for visual place recognition tasks. By extending feature-engineered VLAD into a data-driven learning approach, NetVLAD-based approaches [arandjelovic2016netvlad] have resulted in higher accuracy in terms of matching than the traditional methods. However, extracting 3D spatial geometric features from static images is still challenging. On the other hand, numerous works directly consume point clouds as their inputs.  [uy2018pointnetvlad] employs PointNet [qi2017pointnet] structure and upgrade  [arandjelovic2016netvlad] to directly extract features from a 3d point cloud. To improve the limitation of  [uy2018pointnetvlad] on capturing local geometric structures,  [komorowski2021minkloc3d] proposed to compute more discriminative 3D descriptors based on a sparse voxelized point cloud representation and sparse 3D convolutions.  [du2020dh3d, yang2022fd] designed a Siamese network that jointly learns 3D local feature detection and description directly from raw 3D points. While point clouds offer numerous advantages such as encoded 3D spatial information and viewpoint invariance, it is arguable that integrating 3D features with 2D color features into the descriptors could potentially enhance performance further as this integration exploits the complementary nature of 3D and 2D features, leveraging the strengths of both modalities for more comprehensive scene representation.
@@
Feature Processing: Traditional methods [rublee2011orb, cummins2008fab] use manual-crated features for place recognition. After that CNN-based methods demonstrated better performance in different vision tasks [uy2018pointnetvlad, jin2017learned, gordo2017end] and also place recognition. NetVLAD-based approaches [arandjelovic2016netvlad] show good performance for RGB place recognition, and KPConv-based approaches [Shi2023LiDARBasedPR, ming2022cgis] show promising results in point cloud place recognition. Vision transformers (ViTs) [dosovitskiy2020image] is the next generation of the feature extractor and outperform CNN-based approaches in multiple vision tasks [touvron2021training, berton2022deep, zhu2023r2former]. Recently, a new light-weighted feature extraction method Context of Clusters (CoCs) [cocs], is proposed and demonstrates state-of-the-art performance in different vision tasks but is more computationally effective than ViTs. However, CoCs is only designed for RGB features. In our work, we generalize it to RGB-D point cloud feature extraction.
@@ III-B Architecture of Our PoCo Model
Reducer Blocks: Each reducer block is used to (1) downsample points population from the last layer; (2) estimate relative geometric relationships and aggregate embedding features for reduced points. For points downsampling, we choose the Farthest Point Sampling (FPS) strategy [eldar1997farthest] for its excellent computational efficiency, with the complexity of $O(nk)$, while preserving the structure of the point cloud. Then for each point $p\in\mathbf{P}^{l+1}_{r}$, we apply KNN [kramer2013k] to choose the K nearest points from $\mathbf{P}^{l}_{r}$ as shown in Fig. 3, where $\mathbf{P}^{l}_{r}$ and $\mathbf{P}^{l+1}_{r}$ represent input and output point sets in the reducer block. The feature aggregation from $\mathbf{P}^{l}_{r}$ to $\mathbf{P}^{l+1}_{r}$ is based on PointConvFormer [wu2023pointconvformer] as shown in Eq. 1, where $\mathcal{N}(p)$ represents $p$’s neighbor set. $f_{1,2,3,4}(\cdot)$ are learnable Linear Layers, and $\psi(\cdot)$ is a multi-head attention module calculating the similarity score between $f(p_{k})$ and $f(p)$.
@@
Aggregation: For the point set $\mathbf{P}^{l+1}_{r}$, we downsample it to generate center points $\mathbf{P}^{l+1}_{c}$, where the cardinality $|\mathbf{P}^{l+1}_{r}|=N_{r}>|\mathbf{P}^{l+1}_{c}|=N_{c}$. $N_{c}$ is chosen by FPS, less than half of the $N_{r}$. Each initialized center feature $\mathbf{f}^{\prime}_{c}$ is calculated as the mean vector of its K spatially nearest point features $\{\mathbf{f}_{p}\}$ from $\mathbf{P}^{l+1}_{r}$. As is shown in Fig. 4,in the aggregation stage, we firstly calculate the cosine similarity scores between all the centers and all the points: $s_{c,p}=\sigma(\alpha\cdot cosine(\mathbf{f}_{p},\mathbf{f}^{\prime}_{c})+\beta)$, where $\sigma(\cdot)$ is the sigmoid activation function and $\alpha$ and $\beta$ are trainable parameters. Then point features are aggregated into center points according to previous computed similarities. The center feature $\mathbf{f}^{\prime}_{c}$ is incorporated as an anchor feature for numerical stability as well as further emphasize the locality in Eq. 3
$\mathbf{f}_{c}=\frac{1}{\mathcal{C}}\left(\mathbf{f}^{\prime}_{c}+\sum_{p=1}^{N_{r}}\mathbf{s}_{c,p}\cdot\mathbf{f}_{p}\right),$ (3)
@@
Dispatch: The previous learned similarities are used to dispatch the aggregated center features to their member cluster points via Eq.4 and $h(\cdot)$ indicates Linear Layers:
@@ III-C Training
The circle loss, $\mathcal{L}_{c}$, is defined as Equation 6. $m=0.2$ is a margin to reinforce the optimization and $\delta_{p}=1-m$ and $\delta_{n}=m$. $cl(\cdot)$ is a clamp function, which clamps the value to 0 if it is negative. $\gamma=1$ is a hyperparameter. $s_{p}$ and $s_{n}$ are cosine similarities of the query-positive frame pair and query-negative frame pair, respectively. The circle loss guides the similarities between the query and positive cases to 1 and negative cases to 0.
@@
$\mathcal{L}_{c}=softplus(\log\sum\exp(\gamma a_{n}(s_{n}-\delta_{n}))-$
@@
Triplet loss, $\mathcal{L}_{t}$, is defined as Equation 7. Triplet loss is to maximize the difference between the distances of the query-positive frame pair and query-negative frame pair. The distance between the query and database descriptors is calculated by the L2 distance function, and the circle loss uese similarity is cosine similarity. Those two metrics can be converted by $d(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{2}_{2}=2-2\;cos(\mathbf{x},\mathbf{y})$ for two vectors $\left\{\mathbf{x},\mathbf{y}\right\}$. The margin is set as $m=0.2$.
@@
In the final loss function, we combine the two losses $\mathcal{L}=\alpha\mathcal{L}_{c}+\beta\mathcal{L}_{t}$. In experiments, we find the best weights for these two terms as $\alpha=10$ and $\beta=0.1$ via hyper-parameters tuning. By assigning a higher weight to the circle loss, the model can rapidly converge, facilitating the differentiation between positive and negative examples. Subsequently, circle loss aids in refining the estimation process, resulting in more accurate similarity values that closely align with the boundaries.
@@ IV-A Dataset and Training Settings
Our PoCo model is implemented by Pytorch and trained on 8 Tesla-V100 GPUs. The input points are constrained to 2000 points through voxelization. Besides $\left\{r,g,b\right\}$ and $\left\{x,y,z\right\}$, we also use the normal vector $\left\{n_{x},n_{y},n_{z}\right\}$ of each point to provide additional relative geometric information as in Eq. 2, which is estimated by the plane within a 0.2-meter radius of the point. For the training strategy, we use cosine annealing scheduler [loshchilov2016sgdr] with Adam optimizer [kingma2014adam]. The learning rate is from $10^{-4}$ to $10^{-7}$.
@@ IV-B Implementations of Comparison and Ablation Study
To evaluate the performance of our approach, PoCo, we compare with different SOTA indoor RGB-D and point cloud-based place recognition approaches, CGiS-Net [ming2022cgis], MinkLoc-3D [komorowski2021minkloc3d], NetVLAD [arandjelovic2016netvlad], PointNet-VLAD [uy2018pointnetvlad], Indoor DH3D [yang2022fd]. All these approaches are trained and tested in the two datasets, separately.
@@
Query Frames PoCo w/o Color PoCo Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Query Frames Vanilla-RGB-PR PoCo Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@ V Conclusion, Limitations and Future Work
Acknowledgment: This work was supported in part by ARO Grants W911NF2310046, W911NF2310352 and U.S. Army Cooperative Agreement W911NF2120076